OpenStudy (anonymous):

how to get ride of x when it is in an exponent? for example 6e^2x=18x^2

OpenStudy (anonymous):

\[e^{2x} = 3x^2\] \[\ln (e^{2x}) =\ln (3x^2)\] \[2x \times \ln (e) =2\times \ln (3x)\] \[2x \times 1 =2\times \ln (3x)\] \[x =\ln (3x)\] \[e^{x}= 3x\] thats as far as i know how to simplify this dont know if its possible to isolate x without crazy functions